OpenStudy (anonymous):

The function f(x) = 4(2)x represents the growth of a butterfly population every year in a remote swamp. Jan wants to manipulate the formula to an equivalent form that calculates five times a year, not just once a year. Which function is correct for Jan's purpose, and what is the new growth rate?

OpenStudy (anonymous):

anyone please help?

OpenStudy (anonymous):

@aryandecoolest

OpenStudy (phi):

You should use ^ to show the x is an exponent, as in \[ f(x) = 4\cdot 2^x\]

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

oh, sorry.

OpenStudy (anonymous):

4(2)^x, that?

OpenStudy (phi):

yes. otherwise people might thing you mean 4*2*x (and you would get the wrong answer)

OpenStudy (anonymous):

I'm uncertain what you mean by your phrase "calculates five times per year." What does that mean to you? Calculate ... what?..

OpenStudy (anonymous):

Ok, do you think you can help me or do you want me type the question?

OpenStudy (anonymous):

@aryandecoolest idk, it's a weird written question

OpenStudy (phi):

I think you use \[ (1+ \frac{r}{n})^{nx} \] where r is the original rate and n is now 5

OpenStudy (anonymous):

If, instead of typing f(x)=20(2)x we were to type f(x)=20(2)5x....how do you think the behavior of the original function would be modified?

OpenStudy (anonymous):

i have read this now 5 times and i am also confused as to what it means "five times a year"

OpenStudy (anonymous):

my best guess is that it means \(4\times 2^{\frac{x}{5}}=4\times \sqrt[5]{2}^x\)

OpenStudy (anonymous):

now in this case \(x\) would represent one fifth of a year

OpenStudy (anonymous):

@phi so it's (1 + r/5)^5(x)?

OpenStudy (anonymous):

As much as I'd like to help further, I don't yet see what was done with that '20'. I do see that if we start with f(x) = 20(2)^x, and insert the multiplier 5, we get f(x)=20(32)x...

OpenStudy (anonymous):

the '20' in the original problem statement represents the original number of butterflies. Note that if x=0, 2^x = 2^0 =1 =f(0)=initial value of f(x). Were we to insert that factor of 5, the resulting formula would look like 20(2)5x=20(32)x which would grow a LOT faster....a lot faster than the original formula would, I mean.

OpenStudy (phi):

yes, I think so. you start with 4(2)^x write the 2 as (1+1) 4(1+1)^x now divide the 2nd 1 by 5, and change x to 5x 4(1+1/5)^(5x) that is my best guess as to what they want

OpenStudy (anonymous):

Unfortunately, I share your uncertainty in regard to how to manipulate this last expression so that it resembles one of the four possible answers. ...All I can suggest, at this point, is that you double check to ensure you've copied the problem statement correctly..

OpenStudy (anonymous):

that could be right as well, but it is a different function from the original one

OpenStudy (anonymous):

(x) = 4(1.15)x; growth rate is 5% f(x) = 4(1.15)5x; growth rate is 15% f(x) = 4(2)x; growth rate is 200% f(x) = 4(2)x, growth rate is 5% these are my choice and @phi I said D

OpenStudy (anonymous):

and @aryandecoolest I did copy this right, it's just flvs wants to use trick words and blah blah stuff that don't make any sense

OpenStudy (phi):

my guess would give 4(1+1/5)^(5x) which is the same as 4(1.2)^(5x), however, they came up with 1.15 ?! instead of 1.2

OpenStudy (phi):

but, thinking about it, they want the same growth rate. the original says after 1 year we get 4(2)^1 = 4*2 = 8 if we use 4(1.15)^(5x) after 1 year we get 4(1.15)^5 = 8.045 which is close to 8 in other words, 4(1.15)^(5x) is the correct answer.

OpenStudy (phi):

did they teach you how to do this type of problem? If so, I would like to see what they say

OpenStudy (anonymous):

no they did not

OpenStudy (anonymous):

that was my last question, i'm going to see if it's correct and see what they say.

OpenStudy (anonymous):

thank you

OpenStudy (phi):

btw, the answer is choice B 4(1.15)^5x , 15% rate

OpenStudy (anonymous):

was it correct @morningskye123